Each time Astrid plays a game of chess against her computer, she will win or draw or lose. For each game, the probability of winning is 0.3.The probability of losing is three times the probability she will draw. What is her probability of losing
Answer by richwmiller(17219)

total must equal 1
w=.3
which leaves 1-.3=.7 for lose and draw
l=3d
l+d=.7
3d+d=.7
4d=.7
d=.7/4=0.175
l=3*.7/4=0.525
check
3*0.175=0.525
.3+0.175+0.525=1
ok